Yarra Fabrication estimates that its manufacturing overhead will be $2,240,000 in year 1. It further estimates that direct maPlease explain in detail including the math
Answer #1

Overhead rate = 2240000/1400000 = 160% of direct material used

Applied overhead = 1550000*1.6 = 2480000

Actual overhead = 2400000

Over applied overhead = 80000

No Account and explanation Debit Credit
Manufacturing overhead 80000
Work in process (451000/8200000*80000) 4400
Finished goods (1066000/8200000*80000) 10400
Cost of goods sold 65200